OpenStudy (anonymous):

Evaluate and write your answer in scientific notation:(4 x 102)(15 x 103) A. 60 x 106 B. 6 x 106 C. 6 x 105

OpenStudy (anonymous):

I got it!

OpenStudy (radar):

Assuming it is this:\[(4 x10^{2})(15 x 10^{3})\]

OpenStudy (radar):

\[(4 x 15)(10^{3+2})=60 (10^{5}=6(10^{6})\]